where does the line y=x+30 intersect the parabola y=6x^2
Answer #1

. . The the y=x+30 and the parabola y = 6x² will intersect at that a where 6x² = x+30 it. 682--30=0 ie. x = +1 NE1)2 + 4x6x301004 (-2.154, 27.846) (2.321, 32.321)